# DCoM: Active Learning for All Learners

### v2

@@ Abstract
Deep Active Learning (AL) techniques can be effective in reducing annotation costs for training deep models. However, their effectiveness in low- and high-budget scenarios seems to require different strategies, and achieving optimal results across varying budget scenarios remains a challenge. In this study, we introduce Dynamic Coverage & Margin mix (DCoM), a novel active learning approach designed to bridge this gap. Unlike existing strategies, DCoM dynamically adjusts its strategy, considering the competence of the current model. Through theoretical analysis and empirical evaluations on diverse datasets, including challenging computer vision tasks, we demonstrate DCoM’s ability to overcome the cold start problem and consistently improve results across different budgetary constraints. Thus DCoM achieves state-of-the-art performance in both low- and high-budget regimes.
@@ 1 Introduction
Deep Learning (DL) algorithms require a lot of data to achieve optimal results. In certain scenarios, there is an abundance of unlabeled data but limited capacity for labeling it. In fields such as medical imaging, an invaluable resource — doctors themselves — serves as a costly oracle. Active learning algorithms aim to address this challenge by reducing the burden of labeling and making it more effective. Unlike traditional supervised learning frameworks, Active Learning (AL) can affect the construction of the labeled set, possibly by leveraging knowledge about the current learner. Accordingly, the initial goal of AL is to select $q$ examples to be annotated, where $q$ represents the number of examples that can be sent to the oracle. AL has already demonstrated tangible contributions across various domains such as computer vision tasks [40], NLP [31, 20] and medical imaging [10]. These examples highlight the importance of advancing AL to achieve even greater impact.
@@
The choice of an AL strategy depends on both the learner’s inductive biases and the nature of the problem. Recent research suggests that the optimal active learning strategy varies with the budget size, where budget refers to the size of the training set. When the budget is large, methods based on uncertainty and diversity sampling are most effective. Conversely, when the budget is small, methods centered on typicality and diversity are more appropriate. However, there is no single active learning strategy that is suitable for all budget regimes. Our study aims to address this challenge by dynamic selection of the best examples based on the current state of the learner and the current budget.
@@
As briefly reviewed below, recent research often categorizes active sampling methods according to the budget regime. In contrast, our study takes a different approach by recognizing the need to adapt sample selection methods based on the learner’s competence. We aim to develop a single algorithm that dynamically adjusts its sampling strategy in response to the learner’s evolving abilities. This transition from budget to learner’s competence represents a significant advancement in active learning research, as it allows us to move beyond rigid budget allocations and towards a more adaptive approach that optimizes performance based on the learner’s abilities.
@@
In this work, we introduce a novel method called Dynamic Coverage & Margin mix (DCoM), which dynamically adjusts its selection strategy to deliver optimal results. We begin by examining a simplified theoretical framework (Section 2), where we integrate a typicality test with an uncertainty test based on the budget and its coverage. Inspired by the insights gained from this analysis, we introduce DCoM (Section 3), which combines these characteristics. DCoM aims to offer a flexible solution for any budget by selecting the most appropriate examples that are most suitable to the current budget. We validate DCoM through an extensive empirical study using various vision datasets (Section 4). Our findings indicate that DCoM effectively achieves superior performance across all budget ranges.
@@
Relation to prior art. Over the past years, active learning has been an active area of research [34, 29, 7, 32]. The realm of Active Learning (AL) is segmented into three problem settings: membership query synthesis, stream-based selective sampling, and pool-based active learning [29]. The majority of current AL research adheres to the pool-based AL setting, where samples are drawn from a vast pool of unlabeled data and annotations are solicited from an oracle. In this setting, AL algorithms operate within a given budget ($\mathit{b}$), selecting a subset of unlabeled examples to send to an oracle for labeling. This selection process aims to identify examples that will maximize the model’s performance. This process may be repeated iteratively, gradually increasing budget $b$.
@@
In discussions of pool-based AL, two key characteristics that play a crucial role during the active sampling process are uncertainty and diversity. Some AL algorithms focus solely on uncertainty sampling [23, 30, 4, 43, 18, 12, 27, 37, 11], while others prioritize diversity sampling [28, 19, 39, 15]. Other algorithms combine both uncertainty and diversity in their approaches [38, 2, 36, 21, 5, 1].
@@
Over time, the cold start problem emerged, highlighting the challenges faced by AL algorithms when operating with a small budget. It seems that uncertainty and diversity do not capture this difficulty. Chen et al. [2], for example, describe the cold start problem as a combination of two issues: unbalanced sampling across different classes (which relates to diversity) and unbalanced sampling within each class, where uncertain examples are favored over certain ones at the outset.
@@
In low-budget scenarios, typical examples often offer the best outcomes for the model [2, 39, 15, 14]. This implies that in such contexts, AL algorithms should aim to maximize both certainty and diversity. Conversely, in high-budget scenarios, the emphasis shifts towards maximizing uncertainty while maintaining diversity. [14] proposes a method to select the most suitable approach based on the current state. This method involves comparing multiple active learning strategies during runtime, rather than recommending a single universal approach.
@@
Summary of contribution. (i) Expand a theoretical framework to analyze Active Learning (AL) strategies in embedding spaces. (ii) Introduce DCoM, a multi-budget strategy that significantly outperforms previous methods in low and medium budget scenarios and matches their performance in high budget scenarios. (iii) Initiates a transition from a budget-based approach to the learner’s competence, introducing coverage as a measure designed to predict this competence.
@@ 2 Theoretical framework
The framework of active learning adopted here can be formalized as follows: Let ${\mathbb{X}}$ denote the input domain, ${\mathbb{Y}}$ the target domain, and $\mathcal{L}:{\mathbb{X}}\to{\mathbb{Y}}$ the target learner. Before seeing any labels, $\mathcal{L}$ may be a random feasible hypothesis, or it may be initialized by either transfer learning or self training.
Active learning may involve a single step, or an iterative process with repeated active learning steps. In each step of active a learning, the learner actively seeks labels by choosing a set of unlabeled points as a query set, to be labeled by an oracle/teacher. Subsequently, this set of labeled points is added to the learner’s supervised training set, and the learner is retrained or fine-tuned.
As discussed in the introduction, it has been shown that different active learning strategies are suitable for different budgets $b$. Henceforth, $b$ refers to the number of labeled examples known to the learner at the beginning of an active learning step. With a low budget $b$, strategies that do not rely on the outcome of the learner are most suitable. When $b$ is high, it is beneficial to take into consideration the confidence of the learner when choosing an effective query set. What makes a budget high or low was left vague in previous analysis, as it clearly depends on the specific application and dataset.
To achieve an effective active learning protocol suitable for all learners, irrespective of budget, we aim to devise a universal objective function, whose minimization is used to select the query set. Next, we discuss its envisioned form for selecting the query set at the beginning of each AL step.
We begin our discussion by proposing to replace the notion of budget with the notion of competence: a network is competent if its generalization score is relatively high, and vice versa. This proposal is more formally justified in Section 2.3. We then rephrase the intuition stated above as follows: When the learner’s competence is low, the objective function should prioritize typicality and diversity of selected queries (see discussion in the introduction) irrespective of the learner’s predictions. When the learner’s competence is high, its uncertainty in prediction should be given high priority.
@@
In Section 2.2 we discuss the design of ${\mathcal{O}_{low}}(x)$. In order to tackle the challenge of active learning when the competence of $\mathcal{L}$ is low, we utilize the point coverage framework introduced in [39]. This approach relies on self-supervised data representation, which is blind to the learner’s performance. We refine their analysis and obtain an improved generalization bound for the 1-Nearest Neighbor (1-NN) classification model that depends on the local geometry of the data. Beforehand, in Section 2.1 we discuss the necessary preliminaries required to introduce the coverage approach. In Section 2.3 we discuss our choice of ${S_{\mathcal{L}}}$, and connect it to the notion of budget discussed in previous work.
We select ${\mathcal{O}_{high}}(x)$ to equal one minus the normalized lowest response between the two highest softmax outputs of $\mathcal{L}$ at $x$, or Margin, as this is a common measure of prediction uncertainty.
@@ 2.1 Preliminaries
Notice that $\pi(\delta)$ is monotonically decreasing, as can be readily verified.
In [39] it is shown that the generalization error of the 1-NN classifier $f^{N}$ is bounded as follows
${\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\left(1-P(C({\mathbb{L}},\delta))\right)+(1-\pi(\delta)).$ (2)
Subsequently, an algorithm is proposed that minimizes the first term in (2) by maximizing the coverage probability $P(C({\mathbb{L}},\delta)$, while ignoring the second term that is assumed to be fixed.
@@ 2.2 Refined error bound
Define the indicator random variable $I_{\delta}(x)=\mathbbm{1}_{\{x\in{\mathbb{X}}:B_{\delta}(x)\text{ is pure}\}}$. By definition,
@@
Since the distribution of ${\mathbb{X}}$ is not known apriori, we use the empirical distribution to approximate the expected value of this random variable. In accordance, given a labeled set ${\mathbb{L}}=\left\{x_{i}\right\}_{i=1}^{b}$:
$\begin{split}\hat{\pi}(\delta)&=\hat{{\mathbb{E}}}[\mathbbm{1}_{\{x\in{\mathbb{L}}:B_{\delta}(x)\text{ is pure}\}}]=\frac{1}{b}{\mathbb{E}}[\sum_{i=1}^{b}\mathbbm{1}_{\{B_{\delta}(x_{i})\text{ is pure}\}}]=\frac{1}{b}\sum_{i=1}^{b}{\mathbb{E}}[\mathbbm{1}_{\{B_{\delta}(x_{i})\text{ is pure}\}}]\\ &=\frac{1}{b}\sum_{i=1}^{b}P(B_{\delta}(x_{i})\text{ is pure}).\end{split}$ (3)
@@
${\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\Big{[}1-P(C({\mathbb{L}},\delta))\Big{]}+\Big{[}1-\frac{1}{b}\sum_{i=1}^{b}P(B_{\delta}(x_{i})\text{ is pure})\Big{]}+\varepsilon.$ (4)
@@
where $\varepsilon$ bounds the error of the empirical approximation in (3).
@@
Note that the refined bound in (4) depends on the purity separately at each labeled point in ${\mathbb{L}}$. We further refine this bound by allowing the fixed raidus $\delta$ to be chosen individually, where $\delta_{i}$ denotes the radius of point $x_{i}\in{\mathbb{L}}$. Let $\Delta=[\delta_{i}]_{i=1}^{b}$ denote the list of individual radii corresponding to ${\mathbb{L}}=\{x_{i}\}_{i=1}^{b}$. The cover defined by $\Delta$ is $C({\mathbb{L}},\Delta)=\bigcup_{(x_{i},\delta_{i})\in{\mathbb{L}}\times\Delta}B_{\delta_{i}}(x_{i})$.
We consider the normalized Nearest Neighbor (nNN) classification algorithm, where the nearest neighbor is chosen by solving $f^{N}(x))=\text{argmin}_{x_{i}\in{\mathbb{L}}}\frac{d(x,x_{i})}{\delta_{i}}$. The bound on the error in (4) becomes
@@
${\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\Big{[}1-P(C({\mathbb{L}},\boldsymbol{\Delta}))\Big{]}+\Big{[}1-\frac{1}{b}\sum_{i=1}^{b}P(B_{\delta_{i}}(x_{i})\text{ is pure})\Big{]}+\varepsilon.$ (5)
2.3 Competence score $\boldsymbol{{S_{\mathcal{L}}}}$
@@
In order to obtain a useful competence score [^1] When the labeled set ${\mathbb{L}}$ is large, it may be possible to set aside a validation set to directly estimate the competence of learner $\mathcal{L}$, but this is not feasible when the labeled set ${\mathbb{L}}$ is small., which can be effectively used in objective function (1), we require that it satisfies the following conditions:
  1. (i)
    Lie in the range $[0,1]$.
  2. (ii)
    Use all the data.
  3. (iii)
    Monotonically increase with the competence of learner $\mathcal{L}$.
In accordance, we propose to use the probability of coverage:
${S_{\mathcal{L}}}({\mathbb{L}},\Delta)=\frac{1+e^{-k(1-a)}}{1+e^{-k(P(C({\mathbb{L}},\Delta))-a)}}$ (6)
This function follows a logistic growth curve, with parameters defined as follows: $a\in(0,1)$ is the midpoint and $k\in(0,\infty)$ determines the curve’s steepness.
@@
First, we note that this score satisfies the conditions stated above for the nNN classifier: (i)-(ii), as it is a function in $[0,1]$ that depends on all the data ${\mathbb{X}}$, while (iii) follows from the error bound (5). Below, we prove that it is also monotonically increasing with budget $b$.
@@
More specifically, prior work showed empirically and theoretically that the training budget $b=|{\mathbb{L}}|$ provides a useful clue for the desirable trade-off between ${{\mathcal{O}_{low}}}(x)$ and ${{\mathcal{O}_{high}}}(x)$ in (1), which is captured by score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ defined in (6). We next show that ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ is monotonically increasing with $b$. This implies that definition (6) agrees with empirical evidence in prior work.
@@
${\mathbb{L}}\subseteq{\mathbb{L}}^{\prime}\implies C({\mathbb{L}},\Delta)=\bigcup_{x_{i}\in{\mathbb{L}}}B_{\delta_{i}}(x_{i})\subseteq\left(\bigcup_{x_{i}\in{\mathbb{L}}}B_{\delta_{i}}(x_{i})\right)\cup\left(\bigcup_{x_{i}\in{\mathbb{L}}^{\prime}\setminus{\mathbb{L}}}B_{\delta_{i}}(x_{i})\right)=C({\mathbb{L}}^{\prime},\Delta^{\prime})$
Since probability is monotonic, it follows that $P(C({\mathbb{L}},\Delta))\leq P(C({\mathbb{L}}^{\prime},\Delta^{\prime}))$. From (6), and by the definition of the logistic function using the provided parameters limits, ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ is monotonically increasing in ${\mathbb{L}}$, which implies that ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)\leq{S_{\mathcal{L}}}({\mathbb{L}}^{\prime},\Delta^{\prime})$. ∎

### v3

# DCoM: Competence-Driven Adaptive Active Learning

@@ Abstract
Deep Active Learning techniques can be effective in reducing annotation costs for training deep models. However, their effectiveness in low- and high-budget scenarios seems to require different strategies, and achieving optimal results across varying budget scenarios remains a challenge. In this study, we introduce Dynamic Coverage & Margin mix (DCoM), a novel active learning approach designed to bridge this gap. Unlike existing strategies, DCoM dynamically adjusts its strategy, while taking into account the competence of the current model. Through theoretical analysis and empirical evaluations on diverse datasets, including challenging computer vision tasks, we demonstrate DCoM’s ability to overcome the cold start problem and consistently improve results across different budgetary constraints. Thus DCoM achieves state-of-the-art performance in both low- and high-budget regimes.
@@ 1 Introduction
Deep Learning (DL) algorithms require large amounts of data to achieve optimal results. In some situations, there is an abundance of unlabeled data but limited capacity to label it. In fields such as medical imaging, an invaluable resource — doctors themselves — serves as a costly oracle.
@@
Active learning (AL) algorithms aim to address this challenge by reducing the labeling burden and enhancing its effectiveness (see illustration in Fig. 1). Unlike traditional supervised learning frameworks, AL can influence the construction of the labeled set, potentially by leveraging knowledge about the current learner. Accordingly, the initial goal of AL is to select $q$ examples to be annotated, where $q$ represents the number of examples that can be sent to the oracle. AL has already demonstrated tangible contributions across various domains such as computer vision tasks (e.g. Yuan et al. 2023), NLP (e.g. Siddhant and Lipton 2018; Kaseb and Farouk 2023) and medical imaging (e.g. El-Hasnony et al. 2022). These examples highlight the importance of advancing AL to achieve even greater impact.
Figure 1: Illustration of the AL iterative process: DCoM introduces an additional step to estimate the current learner’s competence, which is then utilized during the active sampling phase.
@@
The choice of an active learning (AL) strategy depends on both the learner’s inductive biases and the nature of the problem. Recent research indicates that the optimal AL strategy varies with the budget size, where the budget refers to the size of the training set. For larger budgets, uncertainty and diversity sampling methods are most effective, while for smaller budgets, strategies focused on typicality and diversity are more suitable. However, no single AL strategy is universally optimal across all budget levels. Our study addresses this challenge by dynamically selecting the best examples based on the learner’s current state and the available budget.
@@
Recent research, as briefly reviewed below, often categorizes active sampling methods based on budget regimes. In contrast, our study takes a different approach by emphasizing the need to adapt sample selection methods according to the learner’s competence. We aim to develop a single algorithm that dynamically adjusts its sampling strategy in response to the learner’s evolving abilities. This shift from a budget-focused to a competence-driven approach represents a significant advancement in active learning research, allowing for a more adaptive strategy that optimizes performance based on measurable learner’s capabilities.
@@
In this work, we introduce a novel method called Dynamic Coverage & Margin Mix (DCoM), which dynamically adjusts its selection strategy by minimizing a new loss function that incorporates a measure of competence. We start by exploring a simplified theoretical framework (Section 2), where we integrate a typicality test with an uncertainty test based on the budget and a notion of coverage. Building on the insights from this analysis, we introduce DCoM (Section 3), which combines these elements to provide a flexible solution for any budget. Thus, DCoM selects the most appropriate examples suited to the current budget. We validate DCoM through an extensive empirical study using various vision datasets (Section 4), demonstrating that DCoM consistently achieves superior performance across all budget ranges.
@@
Relation to prior art. Over the past years, active learning has been an active area of research (Wang et al. 2023; Settles 2009; Cohn, Ghahramani, and Jordan 1996; Tharwat and Schenck 2023; Xie et al. 2023; Schmidt et al. 2024). The realm of Active Learning (AL) is segmented into three problem settings: membership query synthesis, stream-based selective sampling, and pool-based active learning (Settles 2009). Much of the recent AL work follows the pool-based setting, where samples are selected from a large pool of unlabeled data and annotations are obtained from an oracle. In this setting, AL algorithms operate within a given budget ($\mathit{b}$), selecting a subset of unlabeled examples to send to an oracle for labeling. This selection process aims to identify examples that will maximize the model’s performance. This process may be repeated iteratively, gradually increasing budget $b$.
@@
In discussions of pool-based AL, two key characteristics that play a crucial role during the active sampling process are uncertainty and diversity. Some AL algorithms focus solely on uncertainty sampling (Lewis and Gale 1994; Shannon 1948; Cho et al. 2023; Zhao et al. 2021; Houlsby et al. 2011; Gissin and Shalev-Shwartz 2019; Ranganathan et al. 2017; Woo 2022; Gal, Islam, and Ghahramani 2017; Parvaneh et al. 2022; Jung, Kim, and Lee 2023), while others prioritize diversity sampling (Sener and Savarese 2018; Hu, Mac Namee, and Delany 2010; Yehuda et al. 2022; Hacohen, Dekel, and Weinshall 2022). Other algorithms combine both uncertainty and diversity in their approaches (Yang et al. 2015; Chen et al. 2024; Wen, Pizarro, and Williams 2023; Kirsch, Van Amersfoort, and Gal 2019; Chowdhury, Hamerly, and McGarrity 2024; Ash et al. 2019; Kim et al. 2021).
@@
Over time, the cold start problem has emerged, highlighting the difficulties faced by AL algorithms when operating with a small budget. It seems that uncertainty and diversity are ill-suited to address this challenge. Accordingly, Chen et al. (2024) describe the cold start problem as a combination of two issues: unbalanced sampling across different classes (which relates to diversity). and unbalanced sampling within each class, where uncertain examples are favored over certain ones at the outset.
@@
In low-budget scenarios, typical examples often offer the highest benefits for training the model (Hacohen, Dekel, and Weinshall 2022; Yehuda et al. 2022; Chen et al. 2024). This implies that in such contexts, AL algorithms should aim to maximize both certainty and diversity. Conversely, in high-budget scenarios, the emphasis shifts towards maximizing uncertainty while maintaining diversity. While (Hacohen and Weinshall 2024) proposes a method to select the most suitable approach for the current state, the method involves the comparison of multiple active learning strategies at each state, rather than recommending a single universal approach.
@@
Summary of contribution. (i) Expand a theoretical framework to analyze Active Learning (AL) strategies in embedding spaces. (ii) Introduce DCoM, a universal strategy that significantly outperforms previous methods in low and medium budget scenarios, and matches them in high budget scenarios. (iii) Initiate a transition from a budget-based discussion to the learner’s competence, while introducing coverage as a measure designed to predict competence.
@@ 2 Theoretical Framework
Let $\mathbb{X}$} denote the input domain, $\mathbb{Y}$} the target domain, and $\mathcal{L}:{\mathbb{X}}\to{\mathbb{Y}}$ the trained learner. Before seeing any labels, $\mathcal{L}$ may be a random feasible hypothesis, or it may be initialized by either transfer learning or self training. Let $b$ denote the size of the labeled training set, aka budget.
Active learning may involve a single step, or an iterative process with repeated active learning steps. In each AL step, the learner actively seeks labels by choosing a set of unlabeled points as a query set, to be labeled by an oracle/teacher. Subsequently, this set of labeled points is added to the learner’s supervised training set, and the learner is either retrained or fine-tuned.
As discussed in Section 1, it has been shown that different active learning strategies are suited for different budgets $b$. With a low budget $b$, strategies that do not rely on the outcome of the learner are most suitable. When $b$ is high, it is beneficial to consider the confidence of the learner when choosing an effective query set. What makes a certain budget $b$ high or low is left vague in previous work, as it depends on both the hypothesis space (or architecture) and the dataset.
To achieve an effective active learning protocol suitable for all learners, irrespective of budget, we aim to devise a universal objective function, whose minimization is used to select the query set. We begin by proposing to replace the notion of budget with the notion of competence, designed to track the generalization ability of $\mathcal{L}$ when trained with budget $b$. The notion of competence, we hypothesize, is more universal and less dependent on the type of learner and specific dataset (see discussion in Section 2.3).
We then rephrase the intuition stated above (see Section 1) as follows: When the learner’s competence is low, the objective function should prioritize typicality and diversity of selected queries irrespective of the learner’s predictions. When the learner’s competence is high, its uncertainty in prediction should be given high priority.
@@
After the introduction of necessary preliminaries in Section 2.1 we discuss in Section 2.2 the design of ${\mathcal{O}_{low}}(x)$. We begin with the point coverage framework introduced by Yehuda et al. (2022). This approach relies on self-supervised data representation, which is blind to the learner’s performance. We refine their analysis and obtain an improved generalization bound for the Nearest Neighbor classification model that depends on the local geometry of the data.
Next, in Section 2.3 we discuss our choice of $S_{\mathcal{L}}$}, and connect it to the notion of budget discussed in previous work. Finally, we select ${\mathcal{O}_{high}}(x)$ to equal one minus the normalized lowest response between the two highest softmax outputs of $\mathcal{L}$ at $x$, or Margin, as this is a common uncertainty measure.
@@ 2.1 Preliminaries
Note that $\pi(\delta)$ is monotonically decreasing, as can be readily verified.
In (Yehuda et al. 2022) it is shown that the generalization error of the 1-NN classifier $f^{N}$ is bounded as follows
${\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\left(1-P[C({\mathbb{L}},\delta)]\right)+(1-\pi(\delta)).$ (2)
Subsequently, an algorithm is proposed that minimizes the first term in (2) by maximizing the coverage probability $P[C({\mathbb{L}},\delta)]$, while ignoring the second term that is assumed to be fixed.
@@ 2.2 Refined error bound
Define the indicator random variable $I_{\delta}(x)$ as follows:
$I_{\delta}(x)={\mathds{1}_{\{x\in{\mathbb{X}}:B_{\delta}(x)\text{ is pure}\}}}.$
By definition,
@@
Since the distribution of $\mathbb{X}$} is not known apriori, we use the empirical distribution to approximate the expected value of this random variable. Thus, with labeled set ${\mathbb{L}}=\left\{x_{i}\right\}_{i=1}^{b}$
[…]
@@
$\begin{split}{\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\Big[1-P[C({\mathbb{L}},\delta)]\Big]\qquad\qquad\\ +\Big[1-\frac{1}{b}\sum_{i=1}^{b}P[B_{\delta}(x_{i})\text{ is pure}]\Big]+\varepsilon.\end{split}$ (3)
@@
where $\varepsilon\underset{b\to\infty}{\to}0$ bounds the error of the approximation $\hat{\pi}(\delta)$.
@@
Note that the refined bound in (3) depends on the purity separately at each labeled point in $\mathbb{L}$}. Next, we further refine this bound by allowing the fixed radius $\delta$ to be chosen individually, where $\delta_{i}$ denotes the radius of point $x_{i}\in{\mathbb{L}}$. Let $\Delta=[\delta_{i}]_{i=1}^{b}$ denote the list of individual radii corresponding to ${\mathbb{L}}=\{x_{i}\}_{i=1}^{b}$. The cover defined by $\Delta$ is $C({\mathbb{L}},\Delta)=\bigcup_{(x_{i},\delta_{i})\in{\mathbb{L}}\times\Delta}B_{\delta_{i}}(x_{i})$.
With this choice, it is necessary to replace the 1-NN classifier with a suitable variant of nearest neighbor classification - the normalized Nearest Neighbor (1-nNN) classifier. In 1-nNN, distances to point $x_{i}\in{\mathbb{L}}$ are normalized by the corresponding $\delta_{i}$, after which the nearest neighbor $f^{N}(x)=\text{argmin}_{x_{i}\in{\mathbb{L}}}\frac{d(x,x_{i})}{\delta_{i}}$ is computed. This transition preserves the original problem settings, and guarantees that if a labeled point $c\in{\mathbb{L}}$ assigns labels to other points, it also covers them (see Appendix B). Consequently, if a point $x$ is mislabeled, there exists a $c\in{\mathbb{L}}$ that covers it, leading to a reduction in purity. The bound on the error in (3) becomes
@@
$\begin{split}{\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\Big[1-P[C({\mathbb{L}},\Delta)]\Big]\qquad\qquad\\ +\Big[1-\frac{1}{b}\sum_{i=1}^{b}P[B_{\delta_{i}}(x_{i})\text{ is pure}]\Big]+\varepsilon.\end{split}$ (4)
This bound describes a trade-off between probability coverage and purity, and guides the algorithm design in Section 3. Note that in (4), the cover $P[B_{\delta_{i}}(x_{i})\text{ is pure}]$ is refined locally, using $B_{\delta_{i}}$ instead of $B_{\delta}$.
[…]
@@
In order to obtain a useful competence score [^1] When the labeled set $\mathbb{L}$} is large, it may be possible to set aside a validation set to directly estimate the competence of learner $\mathcal{L}$, but this is not feasible when the labeled set $\mathbb{L}$} is small., which can be effectively used in objective function (1), we require that it satisfies the following conditions: (i) Lie in the range $[0,1]$; (ii) Depend on the entire dataset; (iii) Monotonically increase with the competence of learner $\mathcal{L}$. In accordance, we propose to use a constrained Sigmoid function that depends on the probability of coverage:
${S_{\mathcal{L}}}({\mathbb{L}},\Delta)=\frac{1+e^{-k(1-a)}}{1+e^{-k(P[C({\mathbb{L}},\Delta)]-a)}}$ (5)
@@
This score follows the sigmoid curve, where $a$ denotes its midpoint and $k$ the steepness of the curve. In Appendix C, we analyze the choice of these parameters and its impact on the algorithm. Note that this score satisfies the conditions stated above for the nNN classifier: (i)-(ii), as it is a function in $[0,1]$ that depends on $P[C({\mathbb{L}},\Delta)]=\frac{|C({\mathbb{L}},\Delta)|}{|{\mathbb{X}}|}$, while (iii) follows from the error bound (4).
@@
Below, we demonstrate that $S_{\mathcal{L}}$} is monotonically increasing with the budget $b$, which supports the alignment of definition (5) with empirical evidence from prior work. Specifically, Hacohen, Dekel, and Weinshall (2022) showed both empirically and theoretically that the training budget $b=|{\mathbb{L}}|$ serves as a valuable indicator for the relative suitability of ${{\mathcal{O}_{low}}}(x)$ and ${{\mathcal{O}_{high}}}(x)$, which are employed in (1). In our formulation, this relationship informs a trade-off represented by the score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ defined in (5). It remains to be shown that ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ is indeed monotonically increasing with $b$.
@@
$\begin{split}{\mathbb{L}}\subseteq{\mathbb{L}}^{\prime}&\implies C({\mathbb{L}},\Delta)=\bigcup_{x_{i}\in{\mathbb{L}}}B_{\delta_{i}}(x_{i})\\ &\subseteq\left(\bigcup_{x_{i}\in{\mathbb{L}}}B_{\delta_{i}}(x_{i})\right)\cup\left(\bigcup_{x_{i}\in{\mathbb{L}}^{\prime}\setminus{\mathbb{L}}}B_{\delta_{i}}(x_{i})\right)\\ &=C({\mathbb{L}}^{\prime},\Delta^{\prime})\end{split}$
Since probability is monotonic, it follows that $P[C({\mathbb{L}},\Delta)]\leq P[C({\mathbb{L}}^{\prime},\Delta^{\prime})]$. From (5), and by the definition of the logistic function using the provided parameters limits, ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ is monotonically increasing in $\mathbb{L}$}, which implies that ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)\leq{S_{\mathcal{L}}}({\mathbb{L}}^{\prime},\Delta^{\prime})$. ∎